1. A school organized a fundraising activity for its donations to victims of a recent typhoon. It involves selling 1000 raffle tickets. One ticket costs $100. The lone winner of this raffle will receive free tuition worth $30,00. What is your expected net gain if you buy one raffle ticket?




1
Expert's answer
2022-02-28T14:11:26-0500

we define the following:

probability of winning = 1/1000

probability of losing = 999/1000

The gain or loss of winning = $30,00 - $100 = $2900

The gain or loss of losing = $-100


Thus, the expected net gain if you buy one raffle ticket = $ ( (2900 )*(1/1000) )+ ( (-100) *(999/1000) )

Solving the above we get (2.9) - (99.9) = $ -97 which is the required solution.